Consider a $1,600 deposit earning 7 percent interest per year for six years. What is the future value? How much total interest is earned on the original deposit? How much is interest earned on interest?
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

A=$1600*(1.07)^6

=$1600*1.500730352

=$2401.17(Approx)=Future value

Total interest earned=A-P

=(2401.17-1600)=$801.17

interest earned on original deposit=1600*7%*6 years

=$672

interest earned on interest=801.17-672

=$129.17(Approx).